OpenStudy (anonymous):

multiply and reduce the following 5/6 x 3/5 x 1/7

OpenStudy (wolf1728):

= 5 * 3 * 1 / 6 * 5 * 7 The five's can be cancelled and the 3 and 6 can be reduced to 1/2 and we get = 1 * 1 / 2 * 7 = 1/14

OpenStudy (anonymous):

thank you